Add A734 (base 12) and 128B (base 12).
Answer by greenestamps(13200)



Addition in base 12 (A=10; B=11)


      A734
    + 128B
      ----

In base 10, the addition in the rightmost column is 4+11 = 15; in base 12 that is "13".  The 3 is the first digit in the sum; the 1 gets "carried":

        1
      A734
    + 128B
      ----
         3

In base 10, the addition in the next column is 1+3+8 = 12; in base 12 that is "10".  The 0 is the next digit in the sum; the 1 gets carried:

       11
      A734
    + 128B
      ----
        03

In base 10, the addition in the next column is 1+7+2 = 10; in base 12 that is A.  The next digit in the sum is A and there is no carry:

       11
      A734
    + 128B
      ----
       A03

In the last column, the addition in base 10 is 10+1 = 11; in base 12 that is B.  The last digit in the sum is B and we are done.

       11
      A734
    + 128B
      ----
      BA03